OpenStudy (anonymous):

What is the equation for a line that contains the point (0,-3) (4,-1)

OpenStudy (anonymous):

by inspection: y = 1/2 x - 3

OpenStudy (anonymous):

just picture the two points in your head. Then do change in y (2) over change in x (4) = 1/2. And one of the points is already the y-intercept.

OpenStudy (anonymous):

\[y=kx+m\] \[k = {-1-(-3) \over 4-0} = {1 \over 2} \] \[-3 = {1 \over 2} \times 0 + m\] \[m=-3\] \[y={1 \over 2} x -3\] Test: \[-1 = { 1 \over 2} \times 4 - 3 \] QED